If a function g is such that g(-1)=-7 and g(3)=8 then how do i go about finding an equation for g(x) if g(x) is a linear function???????????
Answer by stanbon(75887)

If a function g is such that g(-1)=-7 and g(3)=8 then how do i go about finding an equation for g(x) if g(x) is a linear function???????????
-----
You have 2 points: (-1,-7) and (3,8)
-----
slope = (8--7)/(3--1) = 15/4
-----------------
Form: y = mx + b
8 = (15/4)3 + b
32/4 = 45/4 + b
b = -13/4
-----
Equation:
y = (15/4)x - (13/4)
=========================
Cheers,
Stan H.
=============